# A Million Three-body Binaries Caught by Gaia

### v2

@@ Abstract
Gaia observations have revealed over a million stellar binary candidates within ${\sim}1\,{\rm kpc}$ of the Sun, predominantly characterized by orbital separations ${>}10^{3}\,{\rm AU}$ and eccentricities ${>}0.7$. The prevalence of such wide, eccentric binaries has proven challenging to explain through canonical binary formation channels. However, recent advances in our understanding of three-body binary formation (3BBF)—new binary assembly by the gravitational scattering of three unbound bodies—have shown that 3BBF in star clusters can efficiently generate wide, highly eccentric binaries. We further explore this possibility by constructing a semi-analytic model of the Galactic binary population in the solar neighborhood, originating from 3BBF in star clusters. The model relies on 3BBF scattering experiments to determine how the 3BBF rate and resulting binary properties scale with local stellar density and velocity dispersion. We then model the Galactic star cluster population, incorporating up-to-date prescriptions for the Galaxy’s star formation history as well as the birth properties and internal evolution of its star clusters. Finally, we account for binary destruction induced by perturbations from stellar interactions before cluster escape and and for subsequent changes to binary orbital elements by dynamical interactions in the Galactic field. Without any explicit fine-tuning, our model closely reproduces both the total number of Gaia’s wide binaries and their separation distribution, and qualitatively matches the eccentricity distribution, suggesting that 3BBF may be an important formation channel for these enigmatic systems.
@@ 1 Introduction
Metallicity may provide a further clue to the origin of Gaia’s wide binaries. In particular, the two component members of each such binary typically exhibit near-identical metallicities, implying a co-natal formation channel (J. J. Andrews et al., 2018, 2019; K. Hawkins et al., 2020). In this vein, proposed formation channels for Gaia’s wide binaries include cluster dissolution (M. B. N. Kouwenhoven et al., 2010; N. Moeckel & M. R. Bate, 2010; N. Moeckel & C. J. Clarke, 2011), the dynamical unfolding of triple systems (B. Reipurth & S. Mikkola, 2012), random pairing of objects by thermodynamic fluctuations (J. Peñarrubia, 2019) or primordially (M. Marks & P. Kroupa, 2012; D. Guszejnov et al., 2023; J. P. Farias et al., 2024), dynamical interactions within tidal streams (J. Peñarrubia, 2021), and primordial disk fragmentation (S. Xu et al., 2023). Yet, these formation channels will either produce thermal wide binaries at a consistent rate (e.g., cluster dissolution) or, more rarely, binary populations that are only mildly superthermal (e.g., dynamical unfolding of triples and disk fragmentation). Whether these mechanisms may simultaneously produce extremely wide and highly eccentric binaries frequently enough to match the present-day populations in the solar neighborhood remains unclear, due to a dearth of Galactic theoretical population studies which account for Galactic scale formation rates, Milky Way evolution, diffusion through the disk, and binary field evolution.
@@
There is, however, a dynamical mechanism that readily and preferentially yields wide and highly eccentric binaries—three-body binary formation (3BBF). 3BBF is a triple encounter of single stars, whose outcome is a bound binary and a more energetic single star (P. Mansbach, 1970; S. J. Aarseth & D. C. Heggie, 1976a; A. V. Tutukov, 1978; J. S. Stodolkiewicz, 1986; J. Goodman & P. Hut, 1993; S. F. Portegies Zwart & S. L. W. McMillan, 2000; D. Heggie & P. Hut, 2003; D. Pooley et al., 2003; N. Ivanova et al., 2005; M. Morscher et al., 2015; N. C. Weatherford et al., 2023; D. Atallah et al., 2024; Y. B. Ginat & H. B. Perets, 2024). These new binaries are often referred to as ‘three-body binaries’ (3BBs) and 3BBF is expected to occur frequently in stellar clusters. Binaries may form in clusters by other means, such as primordial binary formation (e.g. F. H. Shu et al. 1987) or tidal capture (A. C. Fabian et al., 1975; W. H. Press & S. A. Teukolsky, 1977), but the fraction of eccentric wide binaries generated by these channels is significantly lower. This work investigates the capacity of 3BBF to generate the wide eccentric binaries observed by Gaia.
@@
Beginning with simulations of isolated scattering between three unbound bodies, we estimate the general 3BBF rate and resulting semi-major axis and eccentricity distributions. We then propagate these distributions in the context of a star cluster environment via a simplified astrophysical model, accounting for the chance of binary escape, disruption en route out of the cluster, the dependence of these probabilities on cluster evolution, the MW’s evolution, and the evolution of binaries deposited in the Galactic field. Our resulting formation rate, a function of nine parameters, predicts the number and orbital element distributions of binaries in the solar neighborhood that have formed via 3BBF in open (or larger) clusters. While 3BBF is automatically and self-consistently included in direct $N$-body simulations of star clusters (e.g. T. S. van Albada 1968; S. J. Aarseth 1969; P. G. Breen & D. C. Heggie 2012a, b; A. Tanikawa 2013; L. Wang et al. 2016; D. Park et al. 2017; J. Kumamoto et al. 2019; M. Arca Sedda et al. 2023), these simulations are computationally prohibitive at this scale. In the spirit of versatility, we opt for a semi-analytic model.
@@
Radius of the interaction volume for 3BBF, normalized by local mean interparticle distance $k\equiv R_{1}/r_{\rm sep}$ A.1
Binary eccentricity $\epsilon$ A.2
Binary semi-major axis (SMA) $a$ A.2
Radial position in the host star cluster, normalized by the cluster’s tidal radius $\tilde{r}\equiv r/r_{t}$ A.3
Time of 3BBF since the cluster’s birth, normalized by the cluter’s total lifetime $\tilde{\tau}\equiv\tau/\tau_{\rm cl}$ A.4
Initial cluster mass $M_{\rm cl0}$ A.4
Galactocentric radius of the cluster’s orbit and intial radius of the escaping 3BB $R_{G}$ A.4/B.1
Galactocentric radius of the binary’s orbit at time $t+\tau$ $R_{G}^{\prime}$ B.1
Time between Milky Way birth and cluster birth $t$ B.1
@@ 2 The Galactic 3BBF Rate: Fundamental Building Blocks
Accurately predicting the contribution of 3BBF in dynamically active star clusters to binaries in the Galactic field requires the careful assembly of a comprehensive set of distribution functions. They describe the properties and evolution of 3BBs and the star clusters that produce them across MW history. When combined, the final equation may be used to estimate the G3R, the total rate of 3BBF within the Galaxy. This section broadly outlines the philosophy of our calculation and the separable components of the G3R, described further in §D.
@@
$\mathcal{G}_{\rm MW}$ $=\frac{dN_{\rm bin}}{d\Omega}=\frac{A_{*}}{A_{\rm R_{G,s}}}[\tilde{\Gamma}_{F}f_{F}P_{\rm esc}P_{\rm ion}f_{\rm CIMF}$ (1)
×$\quad\times\left(N_{\rm ha}f_{\rm ha}SFH_{\rm ha}D_{\rm ha}+N_{\rm la}f_{\rm la}SFH_{\rm la}D_{\rm la}\right)]$
@@
×$\quad\times\,d\tilde{r}\,d\tilde{\tau}\,dk\,\,dM_{\rm cl0}\,dR_{G}\,dR_{G}^{\prime}\,dt\,\,d\epsilon\,da$
@@
Binary field evolution, whether by secular evolution or scattering encounters in the disk, is incorporated for 3BBs that escape their natal clusters and migrate to the present-day solar neighborhood. The two field evolution prescriptions are “phase mixing” (PM, § C.1) and “cumulative scatter” (CS, § C.2) and are denoted by the operators $\mathcal{F_{\rm pm}}$ and $\mathcal{F_{\rm cs}}$, respectively. To obtain a prediction for the distribution of Gaia binaries, we numerically integrate (via the Monte Carlo method) the expression
@@
Volumetric 3BBF Rate $\tilde{\Gamma}_{\rm F}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},k$A.1
Binary semi-major axis & eccentricity distribution $f_{\rm F}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},\epsilon,a,k$A.2
Binary escape probability $P_{\rm esc}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},a$A.5
Binary ionization probability $P_{\rm ion}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},a$A.5
Initial stellar distribution in disk $f_{\rm d}$ $R_{G}$B.1
Star formation history of disk $SFH_{\rm d}$ $R_{G},t$B.1
Stellar diffusion distribution through disk $D_{\rm d}$ $\tilde{\tau},M_{\rm cl0},R_{G},R_{G}^{\prime},t$B.1
Cluster initial mass function $f_{\rm CIMF}$ $M_{\rm cl0},R_{G},t$B.2
Binary survival probability in disk (only if phase mixing) $P_{\rm d,MW}$ $\tilde{\tau},M_{\rm cl0},R_{G},R_{G}^{\prime},t,a$C.1.1
@@
    $P_{\rm esc}$: The fraction of 3BBs, formed with a velocity greater than their host cluster’s local escape speed, assuming an isolated cluster, $P_{\rm esc}$, is calculated from the 3BB velocity distribution in § A.5. The (Maxwellian) 3BB velocity dispersion is the geometric sum of the cluster’s local velocity dispersion and the velocity kick imparted to the binary from the potential energy released during 3BBF. Naturally, the kick is small, especially for (locally) soft/wide binaries, but the escape rate of hard/tight binaries would be substantially underestimated if the kick were neglected.
@@
    $P_{\rm ion}$: Only a small fraction of escaping, wide 3BBs will avoid ionization (i.e., disruption via encounters with other stars) en route out of the cluster. The ionization survival probability of the escaping binary, $P_{\rm ion}$, is calculated in § A.5 based on the classic ionization cross section from P. Hut & J. N. Bahcall (1983). The escaping binary is integrated along a purely radial path and always travels at the local escape speed from an isolated cluster. This assumption is conservative, overestimating $P_{\rm ion}$. [^1] Selecting only binaries with speeds greater than the escape speed from an isolated cluster (rather than the lower escape speed taking into account tidal effects) guarantees the unbound binaries indeed escape promptly, on the crossing timescale, rather than after many additional ($P_{\rm ion}$-boosting) crossings (T. Fukushige & D. C. Heggie, 2000; N. C. Weatherford et al., 2024). Furthermore, larger relative velocities also reduce $P_{\rm ion}$.
  5. 5.
@@
    These three building blocks are tuned with observationally determined scaling constants to separately characterize the thin and thick disks. Each function individually (and all collectively) integrates to unity: $f_{d}$ over all initial Galactocentric radii, $SFH_{d}$ over time and up to the present age of the MW (assumed to be 12 Gyrs), and $D_{d}$ over all Galactocentric radii at time of 3BBF. It is here that we multiply our linearly separable disk distributions by the number of stars in each disk. Thus, integrating over the entire disk parameter space using the disk building blocks yields the total number of stars in each present-day disk.
@@
    $f_{\rm CIMF}$: Our MW model distributes all stars born at time $t$ and location $R_{G}$ into clusters of initial mass $M_{\rm cl0}$ (also born at time $t$) based on the cluster initial mass function (CIMF), $f_{\rm CIMF}$. The CIMF incorporates a Schechter cutoff (P. Schechter, 1976) that is dependent on the local star formation rate at $R_{G}$ in the disk—implicitly found using the MW model described above. It integrates to unity between $20\,M_{\odot}$ and the Schechter upper mass cutoff, $M_{c}$; see § B.2 for details.
  7. 7.
@@
Figure 1: Left: Total number of three-body binaries, $N_{\rm bin}$, we predict are presently within $200\,\rm pc$ of the Sun. We do not account for any observational incompleteness in our models, aside from limiting our comparison radius to $200\,\rm pc$. The black solid line shows the total number of Gaia binaries observed within $200\,\rm pc$ (K. El-Badry et al., 2021), only including the Gaia binary candidates they identified as having a probability ${<}0.1$ of being a chance alignment of two unbound stars. The dashed lines are the totals having removed binaries with average separations ${<}10^{3}\rm AU$. Our models exclude 3BBF under too strong a tidal field within a cluster using the “tidal isolation” parameter, $\eta_{H}$; it sets the maximum allowed tidal acceleration applied by the cluster to the three scattering bodies. Color denotes the field evolution prescription, phase mixing (PM, § C.1) and cumulative scatter (CS, § C.2). The CS prescription destroys more binaries because the destruction and scattering rate of binaries tends to increase following each subsequent encounter and is the unconditionally dominant effect when compared to PM. Right: The average number, $N_{\rm bin}$, of three-body-binaries (3BBs) that escape from a cluster of initial mass $M_{\rm cl0}$ over its entire lifetime; this is largely independent of Galactocentric radius. Note that the right panel does not consider disruption in the MW disk or the subset that migrates to the solar neighborhood. A very small fraction of the initial cluster population forms 3BBs that escape without disruption while en route out of the cluster, with only $1$ to $3$ 3BBs estimated to ever escape a $M_{\rm cl0}\sim 10^{4}M_{\odot}$ cluster over the course of its life.
@@
Figure 3: The predicted number, $N_{\rm bin}$ (top), of 3BB’s within $200\rm pc$ of the Sun and their probability density function, (bottom), as a function of $\langle s\rangle$, projected average binary separation (Equation 4). [^3] Each colored curve is a model realization with a different “tidal isolation” parameter, $\eta_{H}$ (see §3.1). Our most conservative value, $\eta_{H}=0.1$, is the dashed purple line, denoting a local tidal acceleration $0.1\%$ of the average relative acceleration between bodies during three-body scattering. Larger $\eta_{H}$ yields more binaries, but risks breaking the model assumption of isolated three-body scattering. The Gaia binary separations within 200 pc of the Sun are (black curve) taken from K. El-Badry et al. (2021) (with their ${<}0.1$ binary chance-alignment criterion). The left panel is the final distribution when applying “phase mixing” (PM, C.1) while the right panel uses the “cumulative scatter” prescription (CS, C.2). The CS prescription represents the most physically accurate model—it is highly improbable that a wide binary can secularly evolve in isolation over the course of the MW’s entire history. The overall shape of the separation distribution closely matches the observed distribution when $\langle s\rangle>10^{3}\rm AU$. The total number (i.e., from summing over $N_{\rm bin}$) of 3BBs is the number reported in Figure 1.
@@
Figure 5: The eccentricity parameter $\alpha\equiv\mathrm{d}\ln f/\mathrm{d}\ln\epsilon$ (Equation 5), as a function of the projected average binary separation, $\langle s\rangle$. Note that $\alpha=1$ corresponds to the thermal distribution, while $\alpha>1$ and $\alpha<1$ represent superthermal and subthermal distributions, respectively. [^5] Each colored curve denotes the value of $\eta_{\rm H}$ selected to integrate the G3R (Equation D1); $\eta_{\rm H}$ is the maximum size of the spatial three-body interaction region (defined in §3.1) in units of the triple’s Hill radius. Our preferred model, $\eta_{H}=0.10$ (purple), corresponds to a maximum scattering region $10\%$ the local Hill Radius within a cluster. Our most lax condition, $\eta_{H}=0.22$ (light blue), represents tidal force $1\%$ of the average relative acceleration between bodies. The solid black line is the set of $\alpha$ values found by H.-C. Hwang et al. (2022b). In the left panel the colored dashed lines are the “phase-mixed” distributions (§ C.1.2). The right panel features the final eccentricity distributions following the “cumulative scatter” prescription (§ C.2). Phase mixing imparts only a slight “thermalization,” with the resultant wide binary distributions far more eccentric than what is observed. Rather, by considering dynamical encounters in the field, cumulative scattering imparts a dramatic thermalization, resulting in eccentricities in greater observational agreement with H.-C. Hwang et al. (2022b). Despite the simplicity of our toy model and the uncertainties surrounding Gaia-incompleteness, there is good qualitative agreement: wide binaries tend superthermal while tight binaries tend subthermal.

### v3

@@ Abstract
Gaia observations have revealed over a million stellar binary candidates within ${\sim}1\,{\rm kpc}$ of the Sun, predominantly characterized by orbital separations $10^{3}\,{\rm AU}$ and eccentricities >0.7{>}0.7. The prevalence of such wide, eccentric binaries has proven challenging to explain through canonical binary formation channels. However, recent advances in our understanding of three-body binary formation (3BBF)—new binary assembly by the gravitational scattering of three unbound bodies (3UB)—have shown that 3BBF in star clusters can efficiently generate wide, highly eccentric binaries. We further explore this possibility by constructing a semi-analytic model of the Galactic binary population in the solar neighborhood, originating from 3BBF in star clusters and subsequently migrating to the solar neighborhood within a Hubble time. The model relies on 3BBF scattering experiments to determine how the 3BBF rate and resulting binary properties scale with local stellar density, velocity dispersion, and physically-motivated limits to 3UB encounters within a clusters’ tidal field. The Galactic star cluster population is modeled by incorporating up-to-date prescriptions for the Galaxy’s star formation history as well as the birth properties and internal evolution of its star clusters. Finally, we account for binary disruption induced by perturbations from stellar interactions before cluster dissolution and the subsequent changes and disruption of binary orbital elements induced by dynamical interactions in the Galactic field. Our model closely reproduces the total number of Gaia’s wide binaries and the separation and eccentricity distributions, suggesting that 3BBF may be an important formation channel for these enigmatic systems.
@@ 1 Introduction
Metallicity may provide a further clue to the origin of Gaia’s wide binaries. In particular, the two component members of each such binary typically exhibit near-identical metallicities, implying a co-natal formation channel (J. J. Andrews et al., 2018, 2019; K. Hawkins et al., 2020). In this vein, proposed formation channels for Gaia’s wide binaries include cluster dissolution (M. B. N. Kouwenhoven et al., 2010; N. Moeckel & M. R. Bate, 2010; N. Moeckel & C. J. Clarke, 2011), the dynamical unfolding of triple systems (B. Reipurth & S. Mikkola, 2012), random pairing of objects by thermodynamic fluctuations (J. Peñarrubia, 2019) or primordially (M. Marks & P. Kroupa, 2012; D. Guszejnov et al., 2023; J. P. Farias et al., 2024), dynamical interactions within tidal streams (J. Peñarrubia, 2021), and primordial disk fragmentation (S. Xu et al., 2023). Yet, these formation channels will either produce thermal (or sub-thermal) wide binaries at a consistent rate (e.g., random pairing) or, more rarely, binary populations that are only mildly superthermal (e.g., dynamical unfolding of triples and disk fragmentation). Whether these mechanisms may simultaneously produce extremely wide and highly eccentric binaries frequently enough to match the present-day populations in the solar neighborhood remains unclear due to a dearth of theoretical population studies accounting for Galactic scale wide-binary formation rates, Milky Way evolution, diffusion through the disk, and binary field evolution.
@@
There is, however, a dynamical mechanism that readily and preferentially yields wide and highly eccentric binaries—three-body binary formation (3BBF). 3BBF is a triple encounter of initially unbound single stars, whose outcome is a bound binary and a more energetic single star (P. Mansbach, 1970; S. J. Aarseth & D. C. Heggie, 1976; A. V. Tutukov, 1978; J. S. Stodolkiewicz, 1986; J. Goodman & P. Hut, 1993; S. F. Portegies Zwart & S. L. W. McMillan, 2000; D. Heggie & P. Hut, 2003; D. Pooley et al., 2003; N. Ivanova et al., 2005; M. Morscher et al., 2015; N. C. Weatherford et al., 2023; D. Atallah et al., 2024; Y. B. Ginat & H. B. Perets, 2024). These new binaries—often termed ‘three-body binaries’ (3BBs)—are expected to form frequently in stellar clusters. Binaries may also form in clusters by other means, such as primordial binary formation (e.g. F. H. Shu et al. 1987) or tidal capture (A. C. Fabian et al., 1975; W. H. Press & S. A. Teukolsky, 1977), but the fraction of eccentric wide binaries generated by these channels is significantly lower. This work investigates the capacity of 3BBF to generate the wide eccentric binaries observed by Gaia.
@@
Beginning with simulations of isolated scattering between three unbound bodies, we estimate the general 3BBF rate and resulting semi-major axis and eccentricity distributions. We then propagate these distributions in the context of a star cluster environment via a simplified astrophysical model, accounting for the chance of binary disruption until cluster dissolution, the dependence of these probabilities on cluster evolution, the MW’s evolution, and the evolution of binaries deposited in the Galactic field. Our resulting formation rate, a function of nine parameters, predicts the number and orbital element distributions of binaries in the solar neighborhood that have formed via 3BBF in open (or larger) clusters. While 3BBF is automatically and self-consistently included in direct $N$-body simulations of star clusters (e.g. T. S. van Albada 1968; S. J. Aarseth 1969; P. G. Breen & D. C. Heggie 2012a, b; A. Tanikawa 2013; L. Wang et al. 2016; D. Park et al. 2017; J. Kumamoto et al. 2019; M. Arca Sedda et al. 2023), these simulations are computationally prohibitive at this scale. In the spirit of versatility, we opt for a semi-analytic model.
@@
Radius of the interaction volume for 3BBF, normalized by local mean interparticle distance $k\equiv R_{1}/r_{\rm sep}$ § A.1
Binary eccentricity $\epsilon$ § A.2
Binary semi-major axis (SMA) $a$ § A.2
Radial position in the host star cluster, normalized by the cluster’s tidal radius $\tilde{r}\equiv r/r_{t}$ § A.3
Time of 3BBF since the cluster’s birth, normalized by the cluter’s total lifetime $\tilde{\tau}\equiv\tau/\tau_{\rm cl}$ § A.4
Initial cluster mass $M_{\rm cl0}$ § A.4
Galactocentric radius of the cluster’s orbit and intial radius of the escaping 3BB $R_{G}$ § A.4/B.1
Galactocentric radius of the binary’s orbit at time $t+\tau$ $R_{G}^{\prime}$ § B.1
Time between Milky Way birth and cluster birth $t$ § B.1
@@ 2 The Galactic 3BBF Rate: Fundamental Building Blocks
Accurately predicting the contribution of 3BBF in dynamically active star clusters to the population of binaries in the Galactic field requires the careful assembly of a comprehensive set of distribution functions. These distribution functions describe the properties and evolution of 3BBs and the star clusters that produce them across MW history. When combined, the final equation may be used to estimate the G3R, the total rate of 3BBF within the Galaxy. This section broadly outlines the philosophy of our calculation and the separable components of the G3R, described further in §D.
@@
$\mathcal{G}_{\rm MW}$ $=\frac{dN_{\rm bin}}{d\Omega}=\frac{\text{\makebox[0.0pt][l]{\ooalign{\hfil\hphantom{$V$}\hfil\cr\kern 0.80002pt--\hfil\cr}}}V_{*}}{\text{\makebox[0.0pt][l]{\ooalign{\hfil\hphantom{$V$}\hfil\cr\kern 0.80002pt--\hfil\cr}}}V_{\rm R_{G,s}}}[\tilde{\Gamma}_{F}f_{F}P_{\rm no\,enc}f_{\rm CIMF}$ (1)
$\quad\times\left(N_{\rm ha}f_{\rm ha}SFH_{\rm ha}D_{\rm ha}+N_{\rm la}f_{\rm la}SFH_{\rm la}D_{\rm la}\right)]$
@@
$\quad\times\,d\tilde{r}\,d\tilde{\tau}\,dk\,\,dM_{\rm cl0}\,dR_{G}\,dR_{G}^{\prime}\,dt\,\,d\epsilon\,da$
where $N_{\rm bin}$ is the total number of 3BBs and
V
VRG,s
=δl2/(3RG,sh)4.2×104
@@
Binary field evolution, whether by secular evolution or scattering encounters in the disk, is applied to 3BBs that escape their natal clusters and migrate to the present-day solar neighborhood. The two field evolution prescriptions are “phase mixing” (PM, § C.1) and “cumulative scatter” (CS, § C.2) and are denoted by the operators $\mathcal{F_{\rm pm}}$ and $\mathcal{F_{\rm cs}}$, respectively. To obtain a prediction for the distribution of Gaia binaries, we numerically integrate (via the Monte Carlo method) the expression
$N_{\rm bin}=\int\mathcal{F}\left(\mathcal{G}_{\rm MW}\right)d\Omega.$ (2)
@@
Volumetric 3BBF Rate $\tilde{\Gamma}_{\rm F}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},k$ §A.1
Binary semi-major axis & eccentricity distribution $f_{\rm F}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},\epsilon,a,k$ §A.2
Binary in-cluster survival probability $P_{\rm no\,enc}$ $\tilde{r},\tilde{\tau},M_{\rm cl0},R_{G},a$ §A.5
Initial stellar distribution in disk $f_{\rm d}$ $R_{G}$ §B.1
Star formation history of disk $SFH_{\rm d}$ $R_{G},t$ §B.1
Stellar diffusion distribution through disk $D_{\rm d}$ $\tilde{\tau},M_{\rm cl0},R_{G},R_{G}^{\prime},t$ §B.1
Cluster initial mass function $f_{\rm CIMF}$ $M_{\rm cl0},R_{G},t$ §B.2
Binary survival probability in disk (only if phase mixing) $P_{\rm d,MW}$ $\tilde{\tau},M_{\rm cl0},R_{G},R_{G}^{\prime},t,a$ §C.1.1
@@
    $P_{\rm no\,enc}$: Only a small fraction of wide 3BBs will avoid a perturbing encounter or total ionization (i.e., disruption via encounters with other stars) by the time the cluster dissolves. The binary’s probability of surviving without undergoing an encounter, $P_{\rm no\,enc}$, is calculated in § A.5 and is conservatively assumed to be given by the geometric encounter rate as we are only interested in the fraction of binaries unperturbed by any encounters. We do not consider the alternative—where the binary receives a strong kick upon formation, allowing it to escape from the cluster—for this event relies on the exponentially unlikely event that three unbound bodies with velocities at or in excess of the escape velocity (the tail of the distribution function) may meet in a small volume and form soft binaries with a center-of-mass velocity in excess of the local escape velocity. Additionally, the equal-mass 3UB simulations used to generate the semi-analytic fits of § A.2 generate wide binaries with CoM velocities approximately a factor of $\sqrt{2}$ slower than the initial Maxwellian by which the initial velocities were drawn.
@@
    These three building blocks are tuned with observationally determined scaling constants, separately characterizing the thin and thick disks. Each function individually (and all collectively) integrates to unity: $f_{d}$ over all initial Galactocentric radii, $SFH_{d}$ over time and up to the present age of the MW (assumed to be 12 Gyrs), and $D_{d}$ over all Galactocentric radii at time of 3BBF. It is here that we multiply our linearly separable disk distributions by the number of stars in each disk. Thus, integrating over the entire disk parameter space using the disk building blocks yields the total number of stars in each present-day disk.
  5. 5.
    $f_{\rm CIMF}$: Our MW model distributes all stars born at time $t$ and location $R_{G}$ into clusters of initial mass $M_{\rm cl0}$ (also born at time $t$) based on the cluster initial mass function (CIMF), $f_{\rm CIMF}$. The CIMF incorporates a hybrid of the A. Just et al. (2023) CIMF and a Schechter cutoff (P. Schechter, 1976) that is dependent on the local star formation rate at $R_{G}$ in the disk—implicitly found using the MW model described above. It integrates to unity between $50\,M_{\odot}$ and the Schechter upper mass cutoff, $M_{c}$; see § B.2 for details.
@@
Figure 1: Left: Total number of three-body binaries, $N_{\rm bin}$, we predict are presently within $100\,\rm pc$ of the Sun. We do not account for any observational incompleteness in our models, aside from limiting our comparison radius. The black solid line shows the total number of Gaia binaries observed within $100\,\rm pc$ (K. El-Badry et al., 2021), only including the Gaia binary candidates assessed to have a $10\%$ probability of being a chance alignment of two unbound stars. The dashed lines are the totals having removed binaries with average separations $10^{3}\rm AU$. Our models exclude 3BBF under too strong a tidal field within a cluster using the “tidal isolation” parameter, $\eta_{H}$; it sets the maximum allowed tidal acceleration applied by the cluster to the three scattering bodies (Equation 3; see also § A.1). Color denotes the field evolution prescription, phase mixing (PM, § C.1) and cumulative scatter (CS, § C.2). The CS prescription destroys more binaries because the destruction and scattering rate of binaries tends to increase following each subsequent encounter and is the unconditionally dominant effect when compared to PM. Right: The average number, $N_{\rm bin}$, of three-body-binaries (3BBs) that survive without undergoing an encounter in a cluster of initial mass $M_{\rm cl0}$ until cluster dissolution; this is largely independent of Galactocentric radius. Note that the right panel does not consider binary disruption in the MW disk or the subset of binaries that migrate to the solar neighborhood. A very small fraction of the initial cluster population forms 3BBs that survive till dissolution without undergoing an encounter, with ${\lesssim}25\%$ of clusters emitting a single unperturbed 3BB to the field.
@@
Figure 3: The predicted number, $N_{\rm bin}$ (top), of 3BB’s within $100\,\rm pc$ of the Sun and their probability density function (bottom) as a function of $\langle s\rangle$, the projected average binary separation (Equation 4). [^2] Each colored curve is a model realization with a different “tidal isolation” parameter, $\eta_{H}$ (see §3.1). Our most conservative value, $\eta_{H}=0.1$, is the dashed purple line, denoting a local tidal acceleration $0.1\%$ of the average relative acceleration between bodies during three-body scattering. Larger $\eta_{H}$ yields more binaries, but risks breaking the model assumption of isolated three-body scattering. The Gaia binary separations within 100 pc of the Sun are (black curve) taken from K. El-Badry et al. (2021) using their <0.1{<}0.1 binary chance-alignment criterion. The left panel is the final distribution when applying “phase mixing” (PM, C.1) while the right panel uses the “cumulative scatter” prescription (CS, C.2). The CS prescription represents the most physically accurate model. It is highly improbable that a wide binary can secularly evolve in isolation over the course of the MW’s entire history, as is assumed in the PM model. The overall shape of the separation distribution closely matches the observed distribution when $\langle s\rangle>10^{3}\rm AU$. The total number (i.e., from summing over $N_{\rm bin}$) of 3BBs is reported in Figure 1.
@@
Figure 5: The eccentricity parameter $\alpha\equiv\mathrm{d}\ln f/\mathrm{d}\ln\epsilon$ (Equation 5), as a function of the projected average binary separation, $\langle s\rangle$. Note that $\alpha=1$ corresponds to the thermal distribution, while $\alpha>1$ and $\alpha<1$ represent superthermal and subthermal distributions, respectively. [^4] Each colored curve denotes the value of $\eta_{\rm H}$ selected to integrate the G3R (Equation D1); $\eta_{\rm H}$ is the maximum size of the spatial three-body interaction region (defined in §3.1) in units of the triple’s Hill radius. Our preferred model, $\eta_{H}=0.10$ (green), corresponds to a maximum scattering region $10\%$ the local Hill Radius within a cluster. Our most lax condition, $\eta_{H}=0.22$ (light blue), represents a tidal acceleration $1\%$ of the average relative acceleration between bodies. The solid black line is the set of $\alpha$ values found by H.-C. Hwang et al. (2022b). In the left panel the colored dashed lines are the “phase-mixed” distributions (§ C.1.2). The right panel features the final eccentricity distributions following the “cumulative scatter” prescription (§ C.2). Phase mixing imparts only a slight “thermalization,” with the resultant wide binary distributions far more eccentric than what is observed. Rather, by considering dynamical encounters in the field, cumulative scattering imparts a dramatic thermalization, resulting in eccentricities in greater observational agreement with H.-C. Hwang et al. (2022b). Despite the simplicity of our toy model and the uncertainties surrounding Gaia-incompleteness, there is good qualitative agreement: wide binaries tend superthermal while tight binaries tend subthermal.